# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ 4.1 Term languages on graphs
$\llbracket\text{rw}(\bar{u})\rrbracket=[\text{rw}_{u_{i},1},\ldots,\text{rw}_{u_{i},d}]$
@@ 4.2 How powerful is the term language?
Various architectures can be described using this term language. The core idea is always the same: we show that all basic building blocks of the architecture can be captured in the term language and therefore show inductively that every model can be expressed using a term. Let us first note that all linear functions and all commonly used activation functions are Lipschitz continuous, and therefore included in the language.
@@
MPNNs with mean aggregation. Consider an $L$-layer MPNN with mean aggregation, update functions $\small\textsc{Upd}^{(\ell)}$ consisting of an activation function applied to a linear transformation, with mean pooling at the end. First, mean aggregation can be defined using:
$\small\textsc{Mean}_{v}\pi(v)\coloneqq\sum_{y}\pi(y)\star 1$
@@
Now, for each layer $0\leq\ell<L$, we define a term $\tau^{(\ell)}(x)$ which will compute the representation of a node at layer $\ell$ of the MPNN.
  • Let $\tau^{(0)}(x)\coloneqq\mathrm{H}(x)$.
  • For $1\leq\ell<L$, define:
@@
    $\tau^{(\ell+1)}(x)\coloneqq\small\textsc{Upd}^{(\ell)}\left(x,\small\textsc{Mean}_{y\in\mathcal{N}(x)}\tau^{(\ell-1)}(y)\right)$
  • The final representation is computed:
@@
    $\tau\coloneqq\small\textsc{Mean}_{x}\tau^{(L)}(x)$
@@
Using this construction, and letting ${\bm{h}}_{v}^{(\ell)}$ be the representation of the node $v$ in a graph $G$ computed in layer $\ell$ of the MPNN, it is straightforward to verify inductively that:
$\llbracket\tau^{(\ell)}(v)\rrbracket_{G}={\bm{h}}_{v}^{(\ell)}$
@@
The idea is similar for the other architectures. Thus below we only present how the term language captures their respective aggregation functions, which is the fundamental difference between these architectures.
@@
Self-attention and transformers. We can express the self-attention mechanism of transformers using the following aggregator:
$\small\textsc{Att}_{x}\pi(x)\coloneqq\sum_{y}\pi(y)\star\exp(\mathrm{scale}(\pi(x),\pi(y)))$
The function $\mathrm{scale}(\pi(x),\pi(y))$ is a term in the language, since we only need linear transformations with a fixed normalization:
$\frac{\left({\bm{W}}_{Q}\pi(x)\right)^{\top}\left({\bm{W}}_{K}\pi(y)\right)}{\sqrt{d_{k}}}$
for weight matrices ${\bm{W}}_{Q}$ and ${\bm{W}}_{K}$.
With this aggregator we can also capture the GAT architecture.
@@
Graph transformers. To see how graph transformer architectures such as GPS can be expressed, it suffices to note that we can express both self attention layers and MPNN layers with mean aggregation, while the term language is closed under summation. The random walk positional encoding can also be expressed using the rw operator.
@@
Additional architecture features. Because the term language allows the arbitrary combination of graph operations, it can robustly capture many common additional features used in graph learning architectures. For example:
  • A skip connection or residual connection from layer $\ell_{1}$ to layer $\ell_{2}$ can be expressed by including a copy of the term for layer $\ell_{1}$ in the term for the layer $\ell_{2}$ (Xu et al., 2018).
  • Global readout can be captured using a global weighted mean aggregation (Battaglia et al., 2018).
  • Attention conditioned on computed node or node-pair representations can be captured by including the term which computes these representations in the mean weight (Ma et al., 2023).
@@ 5 Convergence theorems
We start out by presenting the convergence theorem for the case of Erdős-Rényi random graphs with node features.
[…]
Let $p\colon\mathbb{N}\to[0,1]$ satisfy one of the following properties.
@@
Concretely, this result shows that for any probabilistic classifier which can be expressed within the term language, when drawing graphs from any of these distributions, eventually the output of the classifier will be the same regardless of input graph, a.a.s. Thus, the only classifiers which can be expressed by such models are those which are asymptotically constant.
Our second result concerns the stochastic block model.
[…]
Our “asymptotically constant” results follow immediately from the above theorems:
@@
We now discuss briefly how the convergence results are proven, starting with the first three cases of Theorem 5.1. We will use that in these cases we have stable behavior in the number of neighbors of nodes. The formalization will be slightly different in each case – see the appendix for details. But in each one neighbourhoods grow at least logarithmically with graph size:
Lemma 5.4.
@@
(Log neighbourhood bound) Let $p\colon\mathbb{N}\to[0,1]$ satisfy density, root growth or logarithmic growth. There is $R>0$ such that for all $\delta,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing graphs from $\mathrm{ER}(n,p(n))$, we have that the proportion of all nodes $u$ such that:
$|\operatorname{\mathcal{N}}(u)|\geq R\log(n)$
is at least $1-\delta$.
@@
While the theorems are about closed terms, naturally we need to prove them inductively on the term language, which requires consideration of terms with free variables. We show that each open term in some sense degenerates to a Lipschitz function almost surely. The only caveat is that we may need to distinguish based on the “type” of the node – for example, nodes $u_{1},u_{2},u_{3}$ that form a triangle may require a different function from nodes that do not. Formally, for node variables $\bar{x}$, an $\bar{x}$ graph type is a conjunction of expressions $E(x_{i},x_{j})$ and their negations. The graph type of tuple $\bar{u}$ in a graph, denoted $\mathrm{GrTp}(\bar{u})$ is the set of all edge relations and their negations that hold between elements of $\bar{u}$. A $(\bar{x},d)$ feature-type controller is a Lipschitz function taking as input pairs consisting of $d$-dimensional real vector and an $\bar{x}$ graph type. A $\wedge$-description (in variables $\bar{x}$) is a conjunction of graph atoms $E(u_{i},u_{j})$. So a $\wedge$-description can be thought of as an incomplete graph type – it does not specify any negated information. The key lemma states that that we can approximate any term of our language with a feature-type controller. In fact we can approximate it relative to any graph type.
Lemma 5.5.
@@
For all terms $\pi(\bar{x})$ over featured graphs with $d$ features, there is a $(\bar{x},d)$ feature-type controller $e_{\pi}$ such that for every $\epsilon,\delta,\theta>0$ and $\wedge$-description $t(\bar{x})$, there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ in the space of graphs of size $n$, out of all the tuples $\bar{u}$ such that $t(\bar{u})$ holds, at least $1-\delta$ satisfy:
$\left\lVert e_{\pi}(H_{G}(\bar{u}),\mathrm{Descr}(\bar{u}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<\epsilon$
@@
This can be seen as a kind of “almost sure quantifier elimination” (thinking of aggregates as quantifiers), in the spirit of Kaila (2003); Keisler & Lotfallah (2009). It is proven by induction on term depth, with the log neighbourhood bound playing a critical role in the induction step for weighted mean.
@@
In the sparse case, we proceed a bit differently. Instead of $\wedge$-descriptions and graph types over $\bar{u}$, which only specify graph relations among the $\bar{u}$, we require descriptions that describe the structure of local neighbourhoods of $\bar{u}$,
@@
In the following we will overload $\operatorname{\mathcal{N}}_{\ell}(\bar{u})$ to indicate the induced subgraph inherited from $G$.
We say two $(k,\ell)$ local neighbourhoods on $\bar{u}$ and $\bar{u}^{\prime}$ are isomorphic if there is a graph isomorphism mapping $\bar{u}$ to $\bar{u}^{\prime}$. A $(k,\ell)$ local isomorphism type is an equivalence class of local neighbourhoods.
For graph $G$ and $\bar{u}$ a $k$-tuple of nodes in $G$, we let $\mathrm{IsoTp}_{\ell}(\bar{u})$ denote the $\ell$ local isomorphism type of $\bar{u}$ in $G$. $\mathsf{IsoTp}_{k,\ell}$ denotes the set of all $(k,\ell)$ local isomorphism types.
For the sparse case, local isomorphism types will play the key role: terms will degenerate to Lipschitz functions of the features and the local isomorphism type. Note that even fixing $k$ and $\ell$, there are infinitely many local isomorphism types, since we do not bound the degree of the nodes.
The combinatorial tool here is that the percentage of local neighbourhoods satisfying a given type converges:
@@
(Local type convergence) Let $p$ satisfy sparsity. For every $k,\ell\in\mathbb{N}$ and $t\in\mathsf{IsoTp}_{k,\ell}$ there is $q_{t}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
||$\left|\frac{|\{u\mid\mathrm{IsoTp}_{\ell}(u)=t\}|}{n}-q_{t}\right|<\epsilon$
Our inductive invariant, analogous to Lemma 5.5 states that every term can be approximated by a Lipschitz function of the features and the local isomorphism type. Formally, for $t\in\mathsf{IsoTp}_{k,\ell}$, a sparse $(t,d)$ controller is a Lipschitz function which takes as input $|t|$-many $d$-vectors from the feature domain and outputs an element of $\mathbb{R}^{d}$
Lemma 5.8.
For every $\pi(x_{1}\ldots x_{k})$, letting $\ell$ be the aggregator depth of $\pi$, for all $t\in\mathsf{IsoTp}_{k,\ell}$ there is a sparse $(t,d)$ controller $e_{\pi}^{t}$ such that for each $\epsilon,\theta>0$, there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ when sampling graphs from $\mathrm{ER}(n,p(n))$ we have that for every $k$-tuple of nodes $\bar{u}$ in the graph such that $\mathrm{IsoTp}_{\ell}(\bar{u})=t$, if $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=\{s_{1},\ldots,s_{|t|}\}$:
Going back to the fact that there are infinitely many local isomorphism types, our controllers a priori to need to be defined for infinitely many $t$. But by local type convergence, eventually all but finitely many types contribute very little. We will use this when we apply the lemma to prove the sparse case of the theorem.
@@ 6 Experimental evaluation
We empirically verify our findings on carefully designed synthetic experiments. The goal of these experiments is to answer the following questions for each model under consideration:
  • Do we empirically observe convergence?
  • What is the impact of the different weighted mean aggregations on the convergence?
  • What is the impact of the (featured) graph distribution on the convergence?
@@
Figure 1: Each plot shows the 5 class probabilities (in different colours) of a single model initialization over the $\mathrm{ER}(n,p(n)=0.1)$ (left column), $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ (middle), and $\mathrm{ER}(n,p(n)=\frac{1}{50n})$ (right column) graph distributions, as we draw increasingly larger graphs.
6.1 Experimental setup
@@
We report experiments for the following architectures: MeanGNN, GAT (Veličković et al., 2018) and GPS+RW (Rampášek et al., 2022). The following setup is carefully designed to eliminate confounding factors:
@@
  • We consider 5 models with the same architecture, each having randomly initialized weights, utilizing a $\operatorname{ReLU}$ non-linearity, and applying a softmax function to their outputs. Each model uses a hidden dimension of $128$, $3$ layers and an output dimension of $5$.
  • We experiment with the following Erdős-Rényi distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ and $\mathrm{ER}(n,p(n)=\frac{1}{50n})$.
@@
  • We experiment with a stochastic block model of 10 communities with equal size, where an edge between nodes within the same community is included with probability $0.7$ and an edge between nodes of different communities is included with probability $0.1$.
@@
  • We draw graphs of sizes up 10,000, where we take 100 samples of each graph size.
  • Node features are independently drawn from $U[0,1]$ and the initial feature dimension is $128$.
  • GPS+RW considers random walks of length up to 5.
@@
The goal of these experiments is to understand the behaviour of the respective models, as we draw larger and larger graphs from the different graph distributions. We independently sample 5 models to ensure this is not a model-specific behaviour.
All our experiments were run on a single NVidia GTX V100 GPU. We made our codebase available online at https://github.com/benfinkelshtein/GNN-Asymptotically-Constant.
Figure 2: Each plot depicts the standard deviation of Euclidean distances between class probabilities and their respective means across various samples of each graph size in GPS+RW, showcasing variations across different graph distributions.
6.2 Empirical results

### v2

@@ 4. For any term composed using $\star$:
$\llbracket\text{rw}(x_{i})(\bar{u})\rrbracket=[\text{rw}_{u_{i},1},\ldots,\text{rw}_{u_{i},d}]$
@@ 4.1 How powerful is the term language?
Various architectures can be described using this term language. The core idea is always the same: we show that all basic building blocks of the architecture can be captured in the term language and applying this inductively yields the desired result. Let us first note that all linear functions and all commonly used activation functions are Lipschitz continuous, and therefore included in the language.
@@
MPNNs with mean aggregation. Consider an $L$-layer MPNN with mean aggregation, update functions $\small\textsc{Upd}^{(\ell)}$ consisting of an activation function applied to a linear transformation, with mean pooling at the end. First, note that mean aggregation can be expressed as:
$\small\textsc{Mean}_{y}\pi(y)\coloneqq\sum_{y}\pi(y)\star 1.$
@@
For each layer $0\leq\ell<L$, we define a term $\tau^{(\ell)}(x)$ which will compute the representation of a node at layer $\ell$ of the MPNN:
  1. Let $\tau^{(0)}(x)\coloneqq\mathrm{H}(x)$. (Initialization)
@@
  2. For $1\leq\ell<L$, define $\tau^{(\ell+1)}(x)\coloneqq\small\textsc{Upd}^{(\ell)}\left(x,\small\textsc{Mean}_{y\in\mathcal{N}(x)}\tau^{(\ell-1)}(y)\right)$. (Layers)
@@
  3. The final graph representation is computed as $\tau\coloneqq\small\textsc{Mean}_{x}\tau^{(L)}(x)$. (Mean pooling)
@@
Using this construction, and letting ${\bm{h}}_{v}^{(\ell)}$ be the representation of the node $v$ in a graph $G$ computed in layer $\ell$ of the MPNN, it is straightforward to verify inductively that $\llbracket\tau^{(\ell)}(v)\rrbracket_{G}={\bm{h}}_{v}^{(\ell)}$.
@@
The idea is similar for the other architectures, where the difference lies in the aggregation functions. Thus below we only present how the term language captures their respective aggregation functions.
@@
Graph transformers. We can express the self-attention mechanism of transformers using the following aggregator:
$\small\textsc{Att}_{y}\pi(y)\coloneqq\sum_{y}\pi(y)\star\exp(\mathrm{scale}(\pi(x),\pi(y)))$
@@
The function $\mathrm{scale}(\pi(x),\pi(y))$ is a term in the language, since scaled dot product attention is a Lipschitz function. To see how graph transformer architectures such as GPS can be expressed, it suffices to note that we can express both self-attention layers and MPNN layers with mean aggregation, since the term language is closed under addition. The random walk positional encoding can also be expressed using the rw operator.
Graph attention networks. The attention mechanism of GAT is local to a node’s neighbours and can be expressed in our term language using similar ideas, except using the local aggregate terms.
@@
Additional architectural features. Because the term language allows the arbitrary combination of graph operations, it can robustly capture many common architectural choices used in graph learning architectures. For example, a skip connection or residual connection from layer $\ell_{1}$ to layer $\ell_{2}$ can be expressed by including a copy of the term for layer $\ell_{1}$ in the term for the layer $\ell_{2}$ [54]. Global readout can be captured using a global mean aggregation [5]. Attention conditioned on computed node or node-pair representations can be captured by including the term which computes these representations in the mean weight [37].
@@
Graph convolutional networks. In Appendix A we show how to extend the term language to incorporate graph convolutional networks (GCNs) [29] by adding a new aggregator.
@@ 5 Convergence theorems
We start by presenting the convergence theorem.
Theorem 5.1.
Consider $(\mu_{n})_{n\in\mathbb{N}}$ sampling a graph $G$ from any of the following models and node features independently from i.i.d. bounded distributions on $d$ features.
  1. 1.
    The Erdős-Rényi distribution $\mathrm{ER}(n,p(n))$ where $p$ satisfies any of the following properties.
@@
  2. 2.
    The Barabási-Albert model $\mathrm{BA}(n,m)$ for any $m\geq 1$.
  3. 3.
    The stochastic block model $\mathrm{SBM}(n_{1}(n),\ldots,n_{m}(n),{\bm{P}})$ where $n_{1},\ldots,n_{m}\colon\mathbb{N}\to\mathbb{N}$ are such that $n_{1}(n)+\cdots+n_{m}(n)=n$ and each $\frac{n_{i}}{n}$ converges, and ${\bm{P}}$ is any symmetric $m\times m$ edge probability matrix.
@@
Concretely, this result shows that for any probabilistic classifier which can be expressed within the term language, when drawing graphs from any of these distributions, eventually the output of the classifier will be the same regardless of input graph, asymptotically almost surely. Thus, the only classifiers which can be expressed by such models are those which are asymptotically constant.
@@
We now discuss briefly how the results are proven. The cases divide into two groups: the denser cases (the first three $\mathrm{ER}$ distributions and the $\mathrm{SBM}$) and the sparser cases (the fourth $\mathrm{ER}$ distribution and the $\mathrm{BA}$ model). Each is proved with a different strategy.
5.1 Overview of the technical constructions for the denser cases
@@
In the denser cases, we first show that we have stable behaviour in the node degrees. The precise formalization will be slightly different in each case, but in each one neighbourhoods grow at least logarithmically with graph size:
Lemma 5.3 (Log neighbourhood bound).
@@
Let $p\colon\mathbb{N}\to[0,1]$ satisfy density, root growth or logarithmic growth. There is $R>0$ such that for all $\delta,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ when drawing graphs from $\mathrm{ER}(n,p(n))$, we have that the proportion of all nodes $u$ such that $|\operatorname{\mathcal{N}}(u)|\geq R\log(n)$ is at least $1-\delta$.
@@
While the theorem is about closed terms, naturally we need to prove it inductively on the term language, which requires consideration of terms with free variables. We show that each open term in some sense degenerates to a Lipschitz function almost surely. The only caveat is that we may need to distinguish based on the “type” of the node – for example, nodes $u_{1},u_{2},u_{3}$ that form a triangle may require a different function from nodes that do not. Formally, for node variables $\bar{x}$, an $\bar{x}$ graph type is a conjunction of expressions $E(x_{i},x_{j})$ and their negations. The graph type of tuple $\bar{u}$ in a graph, denoted $\mathrm{GrTp}(\bar{u})$ is the set of all edge relations and their negations that hold between elements of $\bar{u}$. A $(\bar{x},d)$ feature-type controller is a Lipschitz function taking as input pairs consisting of $d$-dimensional real vector and an $\bar{x}$ graph type.
The key lemma below shows that to each term $\pi$ we can associate a feature-type controller $e_{\pi}$ which captures the asymptotic behaviour of $\pi$, in the sense that with high probability, for most of the tuples $\bar{u}$ the value of $e_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))$ is close to $\llbracket\pi(\bar{u})\rrbracket$.
Lemma 5.4 (Key Lemma).
@@
For all terms $\pi(\bar{x})$ over featured graphs with $d$ features, there is a $(\bar{x},d)$ feature-type controller $e_{\pi}$ such that for every $\epsilon,\delta,\theta>0$, there is $N\in\mathbb{N}$ such that for all $n\geq N$, with probability at least $1-\theta$ in the space of graphs of size $n$, out of all the tuples $\bar{u}$ at least $1-\delta$ satisfy that $\left\lVert e_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<\epsilon$.
@@
This can be seen as a kind of “almost sure quantifier elimination” (thinking of aggregates as quantifiers), in the spirit of Kaila [24], Keisler and Lotfallah [26]. It is proven by induction on term depth, with the log neighbourhood bound playing a critical role in the induction step for weighted mean.
5.2 Overview of the technical constructions for the sparser cases
@@
In the sparser cases, we proceed a bit differently. Instead of graph types over $\bar{u}$, which only specify graph relations among the $\bar{u}$, we require descriptions of local neighbourhoods of $\bar{u}$.
@@
The “types” are now graphs $T$ with $k$ distinguished elements $\bar{w}$, which we call $k$-rooted graphs. The combinatorial tool here is a fact known in the literature as ‘weak local convergence’: the percentage of local neighbourhoods of any given type converges.
Lemma 5.6 (Weak local convergence).
Consider sampling a graph $G$ from either the sparse $\mathrm{ER}$ or $\mathrm{BA}$ distributions. Let $(T,\bar{w})$ be a $k$-rooted graph and take $\ell\in\mathbb{N}$. There is $q_{T}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
||$\left|\frac{|\{\bar{u}\mid\operatorname{\mathcal{N}}_{\ell}(\bar{u})\cong(T,\bar{w})\}|}{n}-q_{T}\right|<\epsilon$
Our inductive invariant, analogous to Lemma 5.4 states that every term can be approximated by a Lipschitz function of the features and the neighbourhood type.
@@
For every $\pi(\bar{x})$, letting $\ell$ be the maximum aggregator nesting depth in $\pi$, for all $k$-rooted graphs $(T,\bar{w})$ there is a Lipschitz function $e_{\pi}^{T}\colon\mathbb{R}^{|T|\cdot d}\to\mathbb{R}^{d}$ such that for each $\epsilon,\theta>0$, there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ in the space of graphs of size $n$, for every $k$-tuple of nodes $\bar{u}$ in the graph such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})\cong(T,\bar{w})$ we have that $\left\lVert e_{\pi}^{T}(H_{G}(\operatorname{\mathcal{N}}_{\ell}(\bar{u})))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<\epsilon$.
@@ 6 Experimental evaluation
We first empirically verify our findings on random graphs and then on a real-world graph to answer the following questions: Q1. Do we empirically observe convergence? Q2. What is the impact of the different weighted mean aggregations on the convergence? Q3. What is the impact of the graph distribution on the convergence? Q4. Can this phenomena arise within large real-world graphs?
All our experiments were run on a single NVidia GTX V100 GPU. We made our codebase available online at %Link␣to␣the␣GitHub␣repositoryhttps://github.com/benfinkelshtein/GNN-Asymptotically-Constant.
@@
Figure 2: Each plot shows the five mean class probabilities (in different colours) with standard deviations of a single model initialization over $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$, and $\mathrm{ER}(n,p(n)=\frac{1}{50n})$, as we draw graphs of increasing size.
@@
Setup. We report experiments for the architectures MeanGNN, GAT [52], and GPS+RW [44] with random walks of length up to 5. Our setup is carefully designed to eliminate confounding factors:
@@
  1. We consider five models with the same architecture, each having randomly initialized weights, utilizing a $\operatorname{ReLU}$ non-linearity, and applying a softmax function to their outputs. Each model uses a hidden dimension of $128$, $3$ layers and an output dimension of $5$.
@@
  2. We experiment with distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$, $\mathrm{ER}(n,p(n)=\frac{1}{50n})$, $\mathrm{BA}(n,m=5)$. We also experiment with an SBM of 10 communities with equal size, where an edge between nodes within the same community is included with probability $0.7$ and an edge between nodes of different communities is included with probability $0.1$.
@@
  3. We draw graphs of sizes up to 10,000, where we take 100 samples of each graph size. Node features are independently drawn from $U[0,1]$ and the initial feature dimension is $128$.
@@
To understand the behaviour of the respective models, we will draw larger and larger graphs from the graph distributions. We use five different models to ensure this is not a model-specific behaviour.
6.1 Empirical results on random graph models